# Quantum geometry from the Moyal product: quantum kinetic equation and non-linear response

### v1

@@ Appendix G Projection operator identities
The trace of the projection operator is not $1$. Rather,
@@
$+\partial^{2}_{\sigma\mu}P_{n}\star\partial_{\alpha}P_{n}+\partial^{2}_{\alpha\mu}P_{n}\star\partial_{\sigma}P_{n}+\partial^{2}_{\alpha\sigma}P_{n}\star\partial_{\mu}P_{n}.$
@@
Also look at Ref. [39] for a further discussion on projection-operator algebra.
Appendix H Projection operator representation of ${\sf H}$ and ${\sf F}$
@@
Appendix I Expressing electric current using $h,f$ (eq. 47)
@@
such that $J_{i}(x)=-e\int_{p}\mathcal{J}_{x_{i}}(x,p)$. Assuming the off-diagonal components of the distribution function, $\tilde{\mathcal{F}}$, are negligible, we can write ${\sf F}=\sum_{n}P_{n}\star\underline{f}_{n}\star P_{n}$ which is the projection operator representation discussed in appendix H. Substituting this into the expression for the phase-space current and using eq. 130, we find
@@
We can now rewrite each of these traces by substituting ${\sf H}=\sum_{n}P_{n}\star h_{n}\star P_{n}$ and using the trace identities of the projection operator listed in appendix G. We find
@@
Appendix J Chiral anomaly effect
@@
Appendix K Deriving $f_{r}$ for the relaxation time approximation (eq. 117)
Here, we derive $f_{r}$ given by eq. 117, the distribution function which $f$ relaxes to under the relaxation time approximation when the Hamiltonian takes the form ${\sf H}={\sf H}_{0}(p)-eV(x)I_{N}$. The relaxation time approximation of the collision integral for the original Liouville-von Neumann equation given by eq. 11 is $-\tau^{-1}({\sf F}-{\sf F}_{r})$ where ${\sf F}_{r}={\sf F}_{\textrm{eq}}=U^{(0)}n_{F}(\varepsilon)U^{(0)\dagger}$. To calculate $f_{r}$, we need to first find an explicit expression for $\tilde{f}_{r}=U^{\dagger}\star{\sf F}_{r}\star U$. To do this, we use $T=\hbar U^{\dagger}\star\partial_{\hbar}U$ which we defined when studying flow equations in appendix D. Recall, $T=Y+\frac{\hbar}{4}\omega^{\alpha\beta}\partial_{\alpha}\Lambda_{\beta}$ where $Y$ is the anti-Hermitian part and the second term is the Hermitian part of $T$. If we expand $T$ in powers of $\hbar$ we find
@@
Appendix L Density-density correlation function

### v2

@@ Appendix H Projection operator identities
The trace of the projection operator is not 11. Rather,
@@
$+\partial^{2}_{\sigma\mu}P_{n}\star\partial_{\alpha}P_{n}+\partial^{2}_{\alpha\mu}P_{n}\star\partial_{\sigma}P_{n}+\partial^{2}_{\alpha\sigma}P_{n}\star\partial_{\mu}P_{n}.\$
@@
Appendix I Projection operator representation of $\sf H$} and $\sf F$}
@@
Appendix J Expressing electric current using $h,f$ (eq. 47)
@@
such that $J_{i}(x)=-e\int_{p}\mathcal{J}_{x_{i}}(x,p)$. Assuming the off-diagonal components of the distribution function, $\tilde{\mathcal{F}}$, are negligible, we can write ${\sf F}=\sum_{n}P_{n}\star\underline{f}_{n}\star P_{n}$ which is the projection operator representation discussed in appendix I. Substituting this into the expression for the phase-space current and using eq. 130, we find
@@
We can now rewrite each of these traces by substituting ${\sf H}=\sum_{n}P_{n}\star h_{n}\star P_{n}$ and using the trace identities of the projection operator listed in appendix H. We find
@@
Appendix K Chiral anomaly effect
@@
Appendix L Deriving $f_{r}$ for the relaxation time approximation (eq. 117)
Here, we derive $f_{r}$ given by eq. 117, the distribution function which $f$ relaxes to under the relaxation time approximation when the Hamiltonian takes the form ${\sf H}={\sf H}_{0}(p)-eV(x)I_{N}$. The relaxation time approximation of the collision integral for the original Liouville-von Neumann equation given by eq. 11 is $\tau^{-1}({\sf F}-{\sf F}_{r})$ where ${\sf F}_{r}={\sf F}_{\textrm{eq}}=U^{(0)}n_{F}(\varepsilon)U^{(0)\dagger}$. To calculate $f_{r}$, we need to first find an explicit expression for $\tilde{f}_{r}=U^{\dagger}\star{\sf F}_{r}\star U$. To do this, we use $T=\hbar U^{\dagger}\star\partial_{\hbar}U$ which we defined when studying flow equations in appendix D. Recall, $T=Y+\frac{\hbar}{4}\omega^{\alpha\beta}\partial_{\alpha}\Lambda_{\beta}$ where $Y$ is the anti-Hermitian part and the second term is the Hermitian part of $T$. If we expand $T$ in powers of $\hbar$ we find
@@
Appendix M Density-density correlation function
@@
Appendix N Comparing to Blount’s mixed representation
In [4, 3, 2], Blount introduced the mixed representation of band theory. In this appendix, we compare our formalism to it.
Consider a separable Hamiltonian given in (89). Using the quantities calculated in Sec.3, we find the band-diagonal but gauge-non-invariant corrections, (147) and (153), as
$\tilde{h}^{(1)}=-\hbar\omega^{\alpha\beta}\partial_{\alpha}\tilde{h}^{(0)}A^{(0)}_{\beta}=e\hbar\partial_{i}VA^{(0)}_{p^{i}}$, (241)
and
$\tilde{h}^{(2)}$ $=-\frac{\hbar}{2}\omega^{\alpha\beta}(\partial_{\alpha}\tilde{h}^{(1)}A_{\beta}^{(0)}+\partial_{\alpha}\tilde{h}^{(0)}A_{\beta}^{(1)})+\frac{1}{2}\tilde{w}^{(2)}$
$=-\frac{\hbar}{2}(\partial_{x^{i}}\tilde{h}^{(1)}A_{p^{i}}^{(0)}+\partial_{x^{i}}\tilde{h}^{(0)}A_{p^{i}}^{(1)})+\frac{\hbar}{2}m^{(1)}_{x^{i}p^{i}}$
$=-\frac{e\hbar^{2}}{2}\partial_{ij}^{2}VA_{p^{i}}^{(0)}A_{p^{j}}^{(0)}-\frac{e^{2}\hbar^{2}}{2}t_{ij}\partial_{i}V\partial_{j}V-\frac{e\hbar^{2}}{2}\partial_{ij}^{2}Vg_{ij}.$ (242)
Gathering above, we have the gauge-non-invariant diagonal Hamiltonian to the second order as
$\tilde{h}$ $=\varepsilon(p)-eV(x)+e\hbar\partial_{i}VA^{(0)}_{p^{i}}-\frac{e\hbar^{2}}{2}\partial_{ij}^{2}VA_{p^{i}}^{(0)}A_{p^{j}}^{(0)}-\frac{e^{2}\hbar^{2}}{2}t_{ij}\partial_{i}V\partial_{j}V-\frac{e\hbar^{2}}{2}\partial_{ij}^{2}Vg_{ij}+\mathcal{O}(\hbar^{3})$
$=\varepsilon(p)-eV\left(x^{i}-\hbar A_{p^{i}}^{(0)}\right)-\frac{e^{2}\hbar^{2}}{2}t_{ij}\partial_{i}V\partial_{j}V-\frac{e\hbar^{2}}{2}\partial_{ij}^{2}Vg_{ij}+\mathcal{O}(\hbar^{3}).$ (243)
We can then define a gauge-invariant diagonal Hamiltonian through
$\underline{h}(x,p)\equiv\tilde{h}(\underline{x},p)$ (244)
where $\underline{x}^{i}\equiv x^{i}+\hbar A_{p^{i}}^{(0)}$, following the notation in [18]. By identifying $x^{i}$ as the quantum operator $\mathrm{i}\frac{\partial}{\partial p^{i}}$, $\underline{h}$ is consistent with the band-diagonal Hamiltonian studied in [4]. In particular, let us consider a Weyl fermion which has a similar quantum geometry as the Dirac fermion in [3] but is non-degenerate. For a right-handed Weyl fermion described by the Hamiltonian $H_{0}(\vec{p})=\vec{p}\cdot\vec{\sigma}$ where $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are the Pauli matrices, the tensors $g_{ij},t_{ij}$ of the negative energy band are given by
$g_{ij}=\frac{1}{4}\partial_{p_{i}}\hat{p}\cdot\partial_{p_{j}}\hat{p}=\frac{1}{4p^{2}}\left(\delta_{ij}-\hat{p}_{i}\hat{p}_{j}\right),$ (245)
where $\hat{p}=\vec{p}/p,p=\lvert{\vec{p}}\rvert$, and $t_{ij}=\frac{1}{p}g_{p_{i}p_{j}}$. Thus, the corrections that involves the gradients of the potential $V$ in $\underline{h}$ are given by
$\Delta\underline{h}=-\frac{e^{2}\hbar^{2}}{8p^{3}}\left(\delta_{ij}-\hat{p}_{i}\hat{p}_{j}\right)\partial_{i}V\partial_{j}V-\frac{e\hbar^{2}}{8p^{2}}\left(\delta_{ij}-\hat{p}_{i}\hat{p}_{j}\right)\partial_{ij}^{2}V+\mathcal{O}(\hbar^{3}).$ (246)
This matches Eq.(12) and Eq.(13) in [3] in the massless limit. Notice that this $\underline{h}$ is not equivalent to our definition of a gauge-invariant diagonal Hamiltonian $h$ in (98). However, as we discussed in the main text, this is merely a choice of representation of the band theory without translational symmetry and will not cause any difference in physical quantities.
Under a uniform magnetic field, the Moyal algebra changes to the magnetic Moyal algebra as described in section 2.9. Assuming the system only depends on the kinetic momentum, i.e. ${\sf H}={\sf H}(\pi_{i})$, we can replace every $\omega^{\alpha\beta}$ with $\omega_{B}^{\pi_{i}\pi_{j}}\equiv-e\mathfrak{F}_{ij}$. Hence, the derivation of the band-diagonal gauge-invariant Hamiltonian follows from (41) and (42). In order to compare with [2], it is instructive to derive the diagonal Hamiltonian using explicitly the gauge field $\Lambda_{p^{i}}$. To this end, we combine (147), (153) and (27) and find that
$h^{(1)}=\frac{i\hbar}{4}\omega^{\alpha\beta}\operatorname{diag}[\{\Lambda^{(0)}_{\alpha}\,,[\Lambda^{(0)}_{\beta}\,,h^{(0)}]\}]=-\frac{e\hbar}{2}\mathfrak{F}_{ij}m_{ij}^{\mathrm{orb}},$ (247)
where $m_{ij}^{\mathrm{orb}}\equiv m_{p^{i}p^{j}}^{(0)}$ is the orbital magnetization. Next, the second order correction is given by
$h^{(2)}$ $=\frac{\hbar}{2}\omega^{\alpha\beta}(\partial_{\alpha}\tilde{h}^{(1)}A_{\beta}^{(0)}+\partial_{\alpha}\tilde{h}^{(0)}A_{\beta}^{(1)})+\frac{1}{2}\tilde{w}^{(2)}$
$\quad+\hbar^{2}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial_{\alpha}\tilde{h}^{(0)}A^{(0)}_{\lambda}\partial_{\sigma}A^{(0)}_{\beta}-\frac{\hbar^{2}}{4}\omega^{\alpha\beta}\omega^{\sigma\lambda}\partial^{2}_{\alpha\sigma}\tilde{h}^{(0)}\left(\operatorname{diag}(\left\{\Lambda^{(0)}_{\beta}\,,\Lambda^{(0)}_{\lambda}\right\})-4A^{(0)}_{\beta}A^{(0)}_{\lambda}\right)$. (248)
Let us temporarily ignore the terms proportional to $Y^{(1)}$ by assuming there is a huge band gap. Then, $h^{(2)}$ simplifies to
$h^{(2)}$ $\approx-(e\hbar)^{2}\mathfrak{u}_{i}\partial_{p_{i}}\left(\frac{1}{2}\mathfrak{F}\cdot m^{\mathrm{orb}}+2v_{j}\mathfrak{u}_{j}\right)-\frac{(e\hbar)^{2}}{2}v_{i}\operatorname{diag}\left(\{\mathfrak{U}_{j},\partial_{p_{j}}\mathfrak{U}_{i}\}-8\mathfrak{u}_{j}\partial_{p_{j}}\mathfrak{u}_{i}\right)$
$\quad\quad+\mathrm{i}\frac{(e\hbar)^{2}}{4}\operatorname{diag}\left(\{\{\mathfrak{U}_{j},\partial_{p_{j}}\mathfrak{U}_{i}\},[\Lambda^{(0)}_{p_{i}},h_{0}]\}\right)-\mathrm{i}\frac{(e\hbar)^{2}}{4}\operatorname{diag}\left(\{\mathfrak{U}_{j},[\Lambda^{(0)}_{p_{j}},\frac{1}{2}\mathfrak{F}\cdot m^{\mathrm{orb}}+2v_{j}\mathfrak{u}_{j}]\}\right)$
$\quad\quad-\frac{(e\hbar)^{2}}{8}\mathfrak{F}_{ij}\operatorname{diag}\left(\partial_{p_{i}}\mathfrak{U}_{k}\partial_{p_{j}}[\Lambda^{(0)}_{p_{k}},h_{0}]+\partial_{p_{i}}[\Lambda^{(0)}_{p_{k}},h_{0}]\partial_{p_{j}}\mathfrak{U}_{k}\right)-\frac{(e\hbar)^{2}}{4}\mathfrak{F}_{ij}\operatorname{diag}\left(\{\mathfrak{U}_{k},\partial_{p_{i}}\Lambda^{(0)}_{p_{k}}v_{j}\}\right)$
$\quad\quad-(e\hbar)^{2}\partial_{p_{i}}v_{j}\left(\operatorname{diag}(\{\mathfrak{U}_{i},\mathfrak{U}_{j}\})-4\mathfrak{u}_{i}\mathfrak{u}_{j}\right)$ (249)
where we defined $\mathfrak{U}_{j}\equiv\frac{1}{2}\mathfrak{F}_{ij}\Lambda^{(0)}_{p_{i}}$ and $\mathfrak{u}_{j}\equiv\operatorname{diag}(\mathfrak{U}_{j})$, and $\mathfrak{F}\cdot m^{\mathrm{orb}}\equiv\mathfrak{F}_{ij}m^{\mathrm{orb}}_{ij}$. These new variables appeared in [2]. First, the magnetization as the first order correction in (247) is consistent with [2]. The second order corrections, however, does not apparently agree with [2]. In fact, we find both the manipulations enigmatic and notations opaque in [2]. For example, if we look at Eq.(I.30) in [2], the term proportional to the velocity $\mathbf{v}_{A}$ seems to trivially vanish since the commutators between diagonal matrices are zero. Our result corresponding to Blount’s $\mathbf{v}_{A}$ would be the second term in the second line of (N) that is, however, nonzero. Moreover, some of the quantities in their Eq.(I.30), like $H$ and $\mathfrak{B}_{\mu}$, are only defined before the unitary rotation, making the comparison rather difficult. Given above, we will not attempt to make a full comparison to Blount’s [2] in this work.